Shift Summary, May 31 3:30-midnight

Douglas A. Jensen (djensen@fnal.gov)
Tue, 1 Jun 1999 00:11:04 -0500 (CDT)

Shift Summary for {Day,Evening,Owl} shift, [date]
On Shift: Doug Jensen

A variety of other committed collaborators were working throughout this
shift.
Jason1, Elizabeth, Arun, Bob H hanging in for a while, Sasha.
Bob T and Ed B came in for beam originally expected between 5-6 pm.
Rick Coleman at MCR, Rick Ford tuning beam from home.

Hours of beam delivered at all intensities: 2.0
Hours of beam delivered at nominal intensity: 2.0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 2

Accounting of lost time:
NM2EU and TeVatron problems until 10 pm, then steady beam. Muon running

No beam this shift due to TeV studies.
- anyway that was the original plan. But changed:
Sometime between noon and 2 pm, the Director decided that
KTeV would get beam as soon as possible. He communicated
this to the division head. The division head informed
Rick Coleman, who informed Bob H ( on shift ). People were
called in to work on the power supply problem.

Activities:

Show up for shift. Status:
Trying to get a run started ( Jason, BobH )
Suspect beam may come - need to monitor status with MCR
Hall ( NM4 ) open
Spooling - mounted a tape ( Jason, BobH )

17:30 UPdate
At 17:15, rain started, and the folks from Main Control arrived
to interlock NM4. Done.
Rick Coleman had come in, reported that magnet repairs proceeding -
klixon change, and water hose mods. Expect to try the magnet around
1800 hrs.
Ed B and Bob T have become part of the discussion and are on their
way in. Hogan informed too. Current ( 1800) plan is to start up
with muons. Hogan can run the 50mv threshold plateaus. Magnet-On
muon run not yet taken.
Discussed with Main Cont. - they are happy to supply whatever intensity
we want. Have requested 1E11 for muon running. I understand we
do not need to write tape for plateaus. Elizabeth stuff for reg gain
matching gets written to disk.

18:30 No beam yet. Quench recovery in progress, People still in NM2.
DC thresholds set to 50 mv
Jaws are closed - to guard against a burst of beam - Rick C in MCR.

20:00 NM2EU ramping, nobody left working at NS7. MCR says supply fixed.
But - TeV problem. Expert on the way in. Est about an hour.


2200 Beam on. Tweeking sweeping magnets to get nice (== as before )
muons.

Hogan doing chamber plateaus, with Sasha Glazov doing analysis,
Elizabeth collecting data for regenerator calib.
These efforts will continue into the next shift.

Note - we discovered in the process of turning on for muons that
we could NOT have the beam stop in and beam turned on. That is,
if the beam stop is IN, the critical device is OFF -> NO BEAM !
Interesting interlock system !!

new shift takes over.

=========================================================================
Runs this shift:
Run 13246 - 13255 for DC plateau. Also Reg Calib.

Tapes brought to FCC this shift: 0

Douglas Jensen DJensen@fnal.gov x8194